xx TestProlongate

TestProlongate

Roland Haas <rhaas@illinois.edu>

August 31 2019

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11 refinedxradius10 Scope: private01  REAL

11 Description: size of the refined region at the center

11 Range 10              01  Default: 0.25
11 (0:*  10 any positive size

11

11 regridxafter 10 Scope: private01  INT

11 Description: when to change the grid

11 Range 10              01  Default: 1
11              10 at initial data time
11 1:*  10 after this iteration

11

7 Interfaces

General Implements: testprolongate

Inherits: carpetxregrid

Grid Variables

7.0.1 PRIVATE GROUPS

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  gf000  10  gf000  01  centering 11 centering={0 0 0}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 dirichletxvalues={1.0e+10}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf001  10  gf001  01  centering 11 centering={0 0 1}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 dirichletxvalues={1.0e+10}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf010  10  gf010  01  centering 11 centering={0 1 0}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 dirichletxvalues={1.0e+10}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf011  10  gf011  01  centering 11 centering={0 1 1}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 dirichletxvalues={1.0e+10}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf100  10  gf100  01  centering 11 centering={1 0 0}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 dirichletxvalues={1.0e+10}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf101  10  gf101  01  centering 11 centering={1 0 1}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 dirichletxvalues={1.0e+10}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  gf110  10  gf110  01  centering 11 centering={1 1 0}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 dirichletxvalues={1.0e+10}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf111  10  gf111  01  centering 11 centering={1 1 1}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 dirichletxvalues={1.0e+10}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf000xerror  10  gf000xerror  01  centering 11 centering={0 0 0}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Error in test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf001xerror  10  gf001xerror  01  centering 11 centering={0 0 1}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Error in test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf010xerror  10  gf010xerror  01  centering 11 centering={0 1 0}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Error in test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf011xerror  10  gf011xerror  01  centering 11 centering={0 1 1}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Error in test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  gf100xerror  10  gf100xerror  01  centering 11 centering={1 0 0}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Error in test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf101xerror  10  gf101xerror  01  centering 11 centering={1 0 1}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Error in test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf110xerror  10  gf110xerror  01  centering 11 centering={1 1 0}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Error in test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf111xerror  10  gf111xerror  01  centering 11 centering={1 1 1}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Error in test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf000xcount  10  gf000xcount  01  centering 11 centering={0 0 0}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Tested grid points
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf001xcount  10  gf001xcount  01  centering 11 centering={0 0 1}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Tested grid points
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  gf010xcount  10  gf010xcount  01  centering 11 centering={0 1 0}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Tested grid points
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf011xcount  10  gf011xcount  01  centering 11 centering={0 1 1}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Tested grid points
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf100xcount  10  gf100xcount  01  centering 11 centering={1 0 0}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Tested grid points
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf101xcount  10  gf101xcount  01  centering 11 centering={1 0 1}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Tested grid points
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf110xcount  10  gf110xcount  01  centering 11 centering={1 1 0}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Tested grid points
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf111xcount  10  gf111xcount  01  centering 11 centering={1 1 1}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Tested grid points
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 gf000xmaxxerror 10 gf000xmaxxerror 01  compact 11 0
11                    10                    01 description 11 Total error in test grid function
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 gf001xmaxxerror 10 gf001xmaxxerror 01  compact 11 0
11                    10                    01 description 11 Total error in test grid function
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 gf010xmaxxerror 10 gf010xmaxxerror 01  compact 11 0
11                    10                    01 description 11 Total error in test grid function
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 gf011xmaxxerror 10 gf011xmaxxerror 01  compact 11 0
11                    10                    01 description 11 Total error in test grid function
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 gf100xmaxxerror 10 gf100xmaxxerror 01  compact 11 0
11                    10                    01 description 11 Total error in test grid function
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 gf101xmaxxerror 10 gf101xmaxxerror 01  compact 11 0
11                    10                    01 description 11 Total error in test grid function
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 gf110xmaxxerror 10 gf110xmaxxerror 01  compact 11 0
11                    10                    01 description 11 Total error in test grid function
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 gf111xmaxxerror 10 gf111xmaxxerror 01  compact 11 0
11                    10                    01 description 11 Total error in test grid function
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 gf000xsumxcount 10 gf000xsumxcount 01  compact 11 0
11                    10                    01 description 11 Number of tested grid points
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 gf001xsumxcount 10 gf001xsumxcount 01  compact 11 0
11                    10                    01 description 11 Number of tested grid points
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 gf010xsumxcount 10 gf010xsumxcount 01  compact 11 0
11                    10                    01 description 11 Number of tested grid points
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 gf011xsumxcount 10 gf011xsumxcount 01  compact 11 0
11                    10                    01 description 11 Number of tested grid points
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 gf100xsumxcount 10 gf100xsumxcount 01  compact 11 0
11                    10                    01 description 11 Number of tested grid points
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 gf101xsumxcount 10 gf101xsumxcount 01  compact 11 0
11                    10                    01 description 11 Number of tested grid points
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 gf110xsumxcount 10 gf110xsumxcount 01  compact 11 0
11                    10                    01 description 11 Number of tested grid points
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 gf111xsumxcount 10 gf111xsumxcount 01  compact 11 0
11                    10                    01 description 11 Number of tested grid points
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11

Uses header: loop.hxx

8 Schedule This section lists all the variables which are assigned storage by thorn CarpetX/TestProlongate. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage

11 Always:  11 x
11 gf000  11 x
11 gf001  11 x
11 gf010xerror  11 x
11 gf011xerror  11 x
11 gf100xerror  11 x
11 gf101xerror  11 x
11 gf110xerror  11 x
11 gf111xerror  11 x
11 gf000xcount  11 x
11 gf001xcount  11 x
11 gf010xcount  11 x
11 gf011xcount  11 x
11 gf010  11 x
11 gf100xcount  11 x
11 gf101xcount  11 x
11 gf110xcount  11 x
11 gf111xcount  11 x
11 gf000xmaxxerror 11 x
11 gf001xmaxxerror 11 x
11 gf010xmaxxerror 11 x
11 gf011xmaxxerror 11 x
11 gf100xmaxxerror 11 x
11 gf101xmaxxerror 11 x
11 gf011  11 x
11 gf110xmaxxerror 11 x
11 gf111xmaxxerror 11 x
11 gf000xsumxcount 11 x
11 gf001xsumxcount 11 x
11 gf010xsumxcount 11 x
11 gf011xsumxcount 11 x
11 gf100xsumxcount 11 x
11 gf101xsumxcount 11 x
11 gf110xsumxcount 11 x
11 gf111xsumxcount 11 x
11 gf100  11 x
11 gf101  11 x
11 gf110  11 x
11 gf111  11 x
11 gf000xerror  11 x
11 gf001xerror  11 x

11 x  11 x

11

Scheduled Functions

CCTKxPOSTINITIAL x testprolongatexregrid x set up test grid x

11 x 11 Language: 11 c
11 x 11 Type:  11 function
11 x 11 Writes: 11 carpetxregrid::regridxerror(interior)

11

CCTKxPOSTSTEP x testprolongatexregrid x set up test grid

11 x 11 Language: 11 c
11 x 11 Type:  11 function
11 x 11 Writes: 11 carpetxregrid::regridxerror(interior)

11

CCTKxINITIAL x testprolongatexsetvalue x set up test data x

11 x 11 Type: 11 group

11

CCTKxEVOL x testprolongatexsetvalue x set up test data x

11 x 11 Type: 11 group

11

TestProlongatexSetValue x testprolongatexset x set up test data x

11 x 11 Language: 11 c
11 x 11 Type:  11 function
11 x 11 Writes: 11 gf000(interior)
11 x 11 x  11 gf001(interior)
11 x 11 x  11 gf010(interior)
11 x 11 x  11 gf011(interior)
11 x 11 x  11 gf100(interior)
11 x 11 x  11 gf101(interior)
11 x 11 x  11 gf110(interior)
11 x 11 x  11 gf111(interior)
11 x 11 x  11 gf000xmaxxerror
11 x 11 x  11 gf001xmaxxerror
11 x 11 x  11 gf010xmaxxerror
11 x 11 x  11 gf011xmaxxerror
11 x 11 x  11 gf100xmaxxerror
11 x 11 x  11 gf101xmaxxerror
11 x 11 x  11 gf110xmaxxerror
11 x 11 x  11 gf111xmaxxerror
11 x 11 x  11 gf000xsumxcount
11 x 11 x  11 gf001xsumxcount
11 x 11 x  11 gf010xsumxcount
11 x 11 x  11 gf011xsumxcount
11 x 11 x  11 gf100xsumxcount
11 x 11 x  11 gf101xsumxcount
11 x 11 x  11 gf110xsumxcount
11 x 11 x  11 gf111xsumxcount

11

TestProlongatexSetValue x testprolongatexsync x sync test data

11 x 11 After:  11 testprolongatexset
11 x 11 Language: 11 c
11 x 11 Sync:  11 gf000
11 x 11 x  11 gf001
11 x 11 x  11 gf010
11 x 11 x  11 gf011
11 x 11 x  11 gf100
11 x 11 x  11 gf101
11 x 11 x  11 gf110
11 x 11 x  11 gf111
11 x 11 Type:  11 function

11

CCTKxPOSTREGRID x testprolongatexsync x sync test data x

11 x 11 Language: 11 c
11 x 11 Sync:  11 gf000
11 x 11 x  11 gf001
11 x 11 x  11 gf010
11 x 11 x  11 gf011
11 x 11 x  11 gf100
11 x 11 x  11 gf101
11 x 11 x  11 gf110
11 x 11 x  11 gf111
11 x 11 Type:  11 function

11

CCTKxPOSTRESTRICT x testprolongatexsync x sync test data x

11 x 11 Language: 11 c
11 x 11 Sync:  11 gf000
11 x 11 x  11 gf001
11 x 11 x  11 gf010
11 x 11 x  11 gf011
11 x 11 x  11 gf100
11 x 11 x  11 gf101
11 x 11 x  11 gf110
11 x 11 x  11 gf111
11 x 11 Type:  11 function

11

CCTKxANALYSIS x testprolongatexcheck x compute error

11 x 11 Language: 11 c
11 x 11 Reads:  11 gf000(everywhere)
11 x 11 x  11 gf001(everywhere)
11 x 11 x  11 gf010(everywhere)
11 x 11 x  11 gf011(everywhere)
11 x 11 x  11 gf100(everywhere)
11 x 11 x  11 gf101(everywhere)
11 x 11 x  11 gf110(everywhere)
11 x 11 x  11 gf111(everywhere)
11 x 11 x  11 gf000xmaxxerror
11 x 11 x  11 gf001xmaxxerror
11 x 11 x  11 gf010xmaxxerror
11 x 11 x  11 gf011xmaxxerror
11 x 11 x  11 gf100xmaxxerror
11 x 11 x  11 gf101xmaxxerror
11 x 11 x  11 gf110xmaxxerror
11 x 11 x  11 gf111xmaxxerror
11 x 11 x  11 gf000xsumxcount
11 x 11 x  11 gf001xsumxcount
11 x 11 x  11 gf010xsumxcount
11 x 11 x  11 gf011xsumxcount
11 x 11 x  11 gf100xsumxcount
11 x 11 x  11 gf101xsumxcount
11 x 11 x  11 gf110xsumxcount
11 x 11 x  11 gf111xsumxcount
11 x 11 Type:  11 function
11 x 11 Writes: 11 gf000xerror(everywhere)
11 x 11 x  11 gf001xerror(everywhere)
11 x 11 x  11 gf010xerror(everywhere)
11 x 11 x  11 gf011xerror(everywhere)
11 x 11 x  11 gf100xerror(everywhere)
11 x 11 x  11 gf101xerror(everywhere)
11 x 11 x  11 gf110xerror(everywhere)
11 x 11 x  11 gf111xerror(everywhere)
11 x 11 x  11 gf000xcount(everywhere)
11 x 11 x  11 gf001xcount(everywhere)
11 x 11 x  11 gf010xcount(everywhere)
11 x 11 x  11 gf011xcount(everywhere)
11 x 11 x  11 gf100xcount(everywhere)
11 x 11 x  11 gf101xcount(everywhere)
11 x 11 x  11 gf110xcount(everywhere)
11 x 11 x  11 gf111xcount(everywhere)
11 x 11 x  11 gf000xmaxxerror
11 x 11 x  11 gf001xmaxxerror
11 x 11 x  11 gf010xmaxxerror
11 x 11 x  11 gf011xmaxxerror
11 x 11 x  11 gf100xmaxxerror
11 x 11 x  11 gf101xmaxxerror
11 x 11 x  11 gf110xmaxxerror
11 x 11 x  11 gf111xmaxxerror
11 x 11 x  11 gf000xsumxcount
11 x 11 x  11 gf001xsumxcount
11 x 11 x  11 gf010xsumxcount
11 x 11 x  11 gf011xsumxcount
11 x 11 x  11 gf100xsumxcount
11 x 11 x  11 gf101xsumxcount
11 x 11 x  11 gf110xsumxcount
11 x 11 x  11 gf111xsumxcount